OpenStudy (anonymous):

A closed box with square base is to be built to house an ant colony. The bottom of the box and all four sides are to be made of material costing 1 dollar/sq ft, and the top is to be constructed of glass costing 5 dollar/sq ft. What are the dimensions of the box of greatest volume that can be constructed for 72 dollars? NOTE: Let denote the length of the side of the base and denote the height of the box

OpenStudy (anonymous):

total cost, if the base has side x and the height is h, is \(1x^2+5x^2+4\times xh=6x^2+5xh=72\) volume is \(V=x^2h\) solve for h in the first equation and substitute in the second to get the volume in terms of x alone

OpenStudy (anonymous):

typo there, should have been \[1x^2+5x^2+4\times xh=6x^2+4xh=72\]

OpenStudy (anonymous):

then \[h=\frac{72-6x^2}{4x}\] and so \[V(x)=x^2\times \frac{72-6x^2}{4x}\] clean this up with some algebra, find the derivative, find the critical points, and that will give a maximimum volume